# Problem Solving Through Human-AI Preference-Based Cooperation

### v2

@@ 5 An exemplary implementation of HAI-Co2
Next, we utilize the code generation module to translate each of the two selected UML candidates into a candidate Python program ( 5 in Figure 3) that will be used for human feedback post-execution  [^5] See https://subha0009.github.io/ExAIC-Interactions/CodeGenLayer.html for a complete round of code generation. Aligned to the goal of co-construction, in this last stage, the user provides their binary judgment on the relative quality of the two generated Python programs along with (optionally) natural language feedback. Such feedback can incorporate the errors found in the program execution (if any), additional requirements, etc. This feedback, along with the summary of the tournament generated by the preference learning module, are together used as a context for the UML generation module to generate a new set of candidate UMLs. This iterative process continues until the user is satisfied with the solution.
Comparison with monolithic LLMs. We compare among GPT-4 Turbo  [^6] See https://subha0009.github.io/ExAIC-Interactions/GPT-4-turbo.html, Meta Llama-3 80B Instruct  [^7] See https://subha0009.github.io/ExAIC-Interactions/Meta-Llama-3-80B-Instruct.html, and HAI-Co2 towards this case study of developing a double pendulum simulation software in Python. We observe multiple immediate improvements with HAI-Co2 over monolithic LLMs that we summarize as follows:
  • Modularity. The incorporation of UML descriptions as interim representation naturally pushes the code generation component in HAI-Co2 towards modular code generation. Single LLMs like GPT-4 Turbo or Llama-3 70B, although able to sometimes follow the instruction of modular code generation, do not respect any defined structure across multiple samples of generation.
  • Ease of iterative modification. The hierarchical abstraction further facilitates seamless iterative modification of the candidate solution across multiple passes of human feedback. Monolithic LLMs tend to introduce unrelated modifications or simply refuse to modify the code despite explicit instructions.
  • Preference adherence. Although our implementation of in-context preference-based search via GPT-4 Turbo is off-the-shelf and quite simple, HAI-Co2 tends to exhibit more transparent and adaptive preference-adherence behavior whereas monolithic LLMs diverge from their initial preference-adherence as the co-construction progresses.
@@
Limitations and further improvements. The immediate improvements we observe are prevalent without any dedicated implementation – neither of the two refinement maps (from natural language to UML and from UML to Python) nor of the preference-based search policy. In this implementation, we do not accommodate direct modification of candidate solutions by the expert as a mode of human feedback. We posit that development along these directions will further improve the quality of co-construction and confirm HAI-Co2’s potential as an effective framework for the class of co-construction problems we aim to address. Modern software development relies on software engineering tools such as type systems, test drivers, static program analysis tools, monitoring and debugging tools and security vulnerability detectors. Realistic software artifacts are complex and their full evaluation by humans without these tools is infeasible. Our implementation of HAI-Co2– not intended as a systematic evaluation of HAI-Co2’s effectiveness in the software domain – will need to be extended with many of the elements that are standard in the DevOps pipeline (see Le et al. (2022); Maninger et al. (2024) for examples of how to integrate such standard tools). We use the LLM’s generative capacity as is, e.g., when it explains why one generated solution is better than another. An interesting research direction would be explanatory interactive learning (Ross et al., 2017; Teso & Kersting, 2019; Friedrich et al., 2023) where more faithful explanations are produced through interactively constraining model explanations.
@@ 6 Conclusion
This paper presents a novel research perspective towards complex problem solving via human-AI co-construction. Our position is that existing generative AI agents require active human participation to successfully construct solutions to complex problems, but cannot effectively serve as reliable partners in human-AI cooperation due to their current limitations. We find evidence for this position in multiple prior research areas across a broad set of domains. Our case study focuses on software generation using GPT-4 Turbo, a strong proprietary LLM, and exemplifies the major drawbacks of current LLMs such as inability to follow human preference, unreliable refinement of complex solution artifacts and limitations to facilitate active human participation. We observed that although day-to-day usage of generative AI tends to adopt a type of human-AI co-construction paradigm in an uninformed manner, the challenges that LLMs face confine such interactions to a much weaker form. As a remedy, we introduce HAI-Co2, a framework that is motivated by the effectiveness of collective human problem solving. HAI-Co2 facilitates a solution construction space with multiple levels of abstractions in which human and AI iteratively refine the candidate solution through search guided by human preference. HAI-Co2 allows active human participation along with versatility in preference expression. After presenting a formalization of HAI-Co2, we discussed the research challenges for this new approach as well as possible future directions for addressing them. Finally, we presented a case study for the application of code generation and noted clear improvements compared to monolithic LLMs.

### v3

@@ 5 An exemplary simulation of HAI-Co2
Next, we utilize the code generation module to translate each of the two selected UML candidates into a candidate Python program that will be used for human feedback post-execution. Aligned to the goal of co-construction, in this last stage, the user provides their binary judgment on the relative quality of the two generated Python programs along with (optionally) natural language feedback. Such feedback can incorporate the errors found in the program execution (if any), additional requirements, etc. This feedback, along with the summary of the tournament generated by the preference learning module, are together used as a context for the next iteration of refinement. This iterative process continues until the user is satisfied with the solution.
Comparison with monolithic LLMs. We perform offline human evaluation to compare HAI-Co2 against a vanilla LLM (in this case, GPT-4 Turbo) in terms of their effectiveness as co-construction partners for expert domain problems. The problem to be solved is to generate code for the double pendulum simulation. Multiple co-construction episodes are generated by specifying different initial preferences and mid-episode preference switching (e.g., choosing different specifications generated by the Specification module; asking for different functionalities in the software; and editing different segments in the specification as well as the codes). Each human evaluator is presented with a pair of co-construction episodes – one using HAI-Co2, one using GPT-4 Turbo – and asked to compare them in terms of the following criteria:
  • Q1.
    Which assistant has better incorporated the initial preferences of the user?
  • Q2.
    Which assistant better adapted to preference switching?
  • Q3.
    Which assistant is more precise in terms of iterative refinement? ‘Precision of modification’ means that the changes are relevant to the request.
  • Q4.
    Which assistant is more complete in terms of iterative refinement? ‘Completeness of modification’ means that all the necessary changes have been made.
  • Q5.
    Overall, which assistant seems more suitable for software-level code generation?
We recruited a total of 14 participants for this evaluation; each is a doctoral student in Natural Language Processing, so that expert preferences in code generation can be understood. On all five criteria, the majority of evaluators rate HAI-Co2 higher than GPT-4 Turbo. 12 (85.7%) evaluators find that HAI-Co2 better captures the initial preferences of the user (Q1). 11 (78.6%) agree that it can adapt better to preference switching (Q2). 11 (78.6%) see HAI-Co2 as superior on precision (Q3) and completeness of modification (Q4). 12 (85.7%) suggest HAI-Co2 is better suited for software development (Q5). We present detailed justifications for the responses provided by the evaluators in Appendix B.
@@
Limitations and further improvements. The immediate improvements we observe are prevalent without any dedicated implementation – neither of the two refinement maps (from natural language to UML and from UML to Python) nor of the preference-based search policy. In this implementation, we do not equip the co-construction space with explicit neighborhood structures. We posit that development along these directions will further improve the quality of co-construction and confirm HAI-Co2’s potential as an effective framework for the class of co-construction problems we aim to address. Modern software development relies on software engineering tools such as type systems, test drivers, static program analysis tools, monitoring and debugging tools and security vulnerability detectors. Realistic software artifacts are complex, and their full evaluation by humans without these tools is infeasible. Our implementation of HAI-Co2– not intended as a systematic evaluation of HAI-Co2’s effectiveness in the software domain – will need to be extended with many of the elements that are standard in the DevOps pipeline (see Le et al. (2022); Maninger et al. (2024) for examples of how to integrate such standard tools). We use the LLM’s generative capacity as is, e.g., when it explains why one generated solution is better than another. An interesting research direction would be explanatory interactive learning (Ross et al., 2017; Teso & Kersting, 2019; Friedrich et al., 2023), where more faithful explanations are produced through interactively constraining model explanations. The search strategy can be further refined by implementing reinforcement learning from execution feedback (Gehring et al., 2024; Liu et al., 2023; Dutta et al., 2024).
@@ 6 Conclusion
This paper presents a novel research perspective towards complex problem solving via human-AI co-construction. Our position is that existing generative AI agents require active human participation to successfully construct solutions to complex problems, but cannot effectively serve as reliable partners in human-AI cooperation due to their current limitations. We find evidence for this position in multiple prior research areas across a broad set of domains. Our case study focuses on software generation using GPT-4 Turbo, a strong proprietary LLM, and exemplifies the major drawbacks of current LLMs such as inability to follow human preferences, unreliable refinement of complex solution artifacts and limitations to facilitate active human participation. We observed that although day-to-day usage of generative AI tends to adopt a type of human-AI co-construction paradigm in an uninformed manner, the challenges that LLMs face confine such interactions to a much weaker form. As a remedy, we introduce HAI-Co2, a framework that is motivated by the effectiveness of collective human problem solving. HAI-Co2 facilitates a solution construction space with multiple levels of abstractions, in which human and AI iteratively refine the candidate solution through search guided by human preference. HAI-Co2 allows active human participation along with versatility in preference expression. After presenting a formalization of HAI-Co2, we discussed the research challenges for this new approach as well as possible future directions for addressing them. Finally, we presented a case study for the application of code generation and noted clear improvements compared to monolithic LLMs.
Appendix A Possible implementation in Related Work Generation
Figure 4: A conceptual application of HAI-Co2 for the problem of generation of a related work section of a research paper. The user provides the title and the introduction of the research paper for which the related work section is to be co-constructed (i.e., written). Three abstraction hierarchies are envisioned. The Literature space consists of lists of papers. The Semantic graph space depicts the papers, their findings, and their interrelations using a directed graph. The Related work space contains related work sections (written in natural language using a writing style appropriate for this genre) that describe the semantic graph. The Paper finder module lists down the papers that should be incorporated into the related work. Neighborhood structure is imposed upon this space using semantic similarity and hops over the citation network. The Preference decoding module keeps track of the user’s preferences. Upon deciding on a suitable list of papers, the Semantic graph module translates them into a semantic graph. Finally, the Related Work (RW) generation module writes related work sections based on the semantic graph.
To showcase the applicability of HAI-Co2 for expert domains other than software engineering, we present a workflow for the problem of related work generation in Figure 4. Note that this is not an actual implementation, rather a proposal on how HAI-Co2 can be adapted for this problem. Similar to our case study on software engineering, a solution construction space with three levels of hierarchy is defined. The highest level of abstraction (Literature space) is the space of lists of relevant papers; each point (represented as a list of papers) is intended to capture the literature relevant to the research paper (in this case "Problem solving through Human-AI Preference-Based Co-operation"). Based on the user’s problem specification, the Paper finder module (which can be an LLM-web search hybrid) lists the possible papers that are relevant to the research paper. This is a classical search problem. Next, these papers are used to construct a semantic graph that relates different papers according to their domain of focus, methodology, findings, prescriptions, etc. Such a graph is inherently heterogeneous. Multiple semantic graphs can be generated from a given list of papers. One can define a neighborhood over the space of these semantic graphs via edit distance. The search strategy, in this case, can again be realized through a tournament (as in the software engineering domain presented in Section 5) or through another mechanism (e.g., a specialized module that evaluates semantic graphs based on their graph-theoretical properties). Finally, the RW (related work) generation module translates these semantic graphs into Related work sections. Locally valid neighborhood structures can be constructed using the neural representations of textual differences. The Preference decoder module extracts the preferences expressed by the user to guide the search in different spaces. In this example, one can define refinement maps between the different abstraction levels in a straightforward manner: The semantic graph can be mapped to the list of papers directly as the former has nodes that are members of the latter. Similarly, each pair of nodes and their connecting edge in the semantic graph can be translated to a sentence in the related work section in the Related work space.
Appendix B Human evaluation
We provide each evaluator with a pair of interactions between i) a user and GPT-4 Turbo  [^7] Example interaction can be found at https://subha0009.github.io/ExAIC-Interactions-10A2/Assistant2.html and ii) a user and HAI-Co2  [^8] Example interaction can be found at https://subha0009.github.io/ExAIC-Interactions-10A2/Assistant1.html. We do not collect any personal information from the evaluators. Detailed responses are available here: https://subha0009.github.io/ExAIC-Interactions-10A2/FormResponses.html.
Figure 5: Interface used for human evaluation.